OpenStudy (anonymous):

For the polynomial 6xy2 – 5x2y? + 9x2 to be a trinomial with a degree of 3 after it has been fully simplified, the missing exponent of the y-term must be ?

OpenStudy (anonymous):

Example 1. Factor 30. Solution. 30 = 2· 15 = 2· 3· 5 If we begin 30 = 5· 6, we still obtain -- apart from the order -- 5· 2· 3. Problem 1. Factor 50. To see the answer, pass your mouse over the colored area. To cover the answer again, click "Refresh" ("Reload"). Do the problem yourself first! 50 = 2· 25 = 2· 5· 5 Factoring, then, is the reverse of multiplying. When we multiply, we write 2(a + b) = 2a + 2b. But if we switch sides and write 2a + 2b = 2(a + b), then we have factored 2a + 2b as the product 2(a + b). In this sum 2a + 2b, 2 is a common factor of each term. It is a factor of 2a, and it is a factor of 2b. This Lesson is concerned exclusively with recognizing common factors. Problem 2. Factor 3x − 3y. 3x − 3y = 3(x − y) Problem 3. Rewrite each of the following as the product of 2x and another factor. For example, 10x3 = 2x· 5x2. Rule 1 of exponents. a) 8x = 2x· 4 b) 6ax = 2x· 3a c) 2x2 = 2x· x d) 2x3 = 2x· x2 e) 4x10 = 2x· 2x9 f) 6x5 = 2x· 3x4 g) 2ax6 = 2x· ax5 h) 2x = 2x· 1

Mercury:

... no idea what the previous guy is on about... Anyway, asked & answered here: https://questioncove.com/study#/updates/5f64be185ca306717f18140